Find the solution set of these simultaneous equations.
2x + 3y = 10 and 3x -2y = 2

Answer by richwmiller(17219)

there are two equations and therefore are two simultaneous equations
2x + 3y = 10 multiply by 2
3x -2y = 2 multiply by 3
4x + 6y = 20
9x -6y = 6
add
13x=26
x=2
plug in 2 for x and solve for y
enjoy